PLEASE HELP ASAP!
Use the formula A = Pe^rt to find the total amount of money accumulated at the end of the indicated time period by compounding continuously.
Round your answer to the nearest cent
$950 for 4 years at 7%

Answer by nerdybill(6962)

Use the formula A = Pe^rt to find the total amount of money accumulated at the end of the indicated time period by compounding continuously.
Round your answer to the nearest cent
$950 for 4 years at 7%
.
simply plug in the given values:
A = Pe^(rt)
A = 950e^(.07*4)
A = 950e^(.28)
A = 950(1.32313)
A = $1256.97